OpenStudy (anonymous):

how do you solve cos4x+cos2x=0?

OpenStudy (solomonzelman):

\[Let~~~~~Cos^2x=a\]

OpenStudy (solomonzelman):

if you meant \[\cos^4x+\cos^2x=0\]

OpenStudy (anonymous):

no i meant like the double angle...

OpenStudy (anonymous):

as it is written

OpenStudy (solomonzelman):

\[\cos(4x)+\cos(2x)=0\]

OpenStudy (anonymous):

yes :)

OpenStudy (jdoe0001):

http://web.mit.edu/wwmath/trig/eq7.gif

OpenStudy (solomonzelman):

\[\cos(2x+2x)+\cos(2x)=0\]\[use:~~~~\cos(A+B)=\cos(A+B)=\cos A \cos B - \sin A \sin B\]

OpenStudy (solomonzelman):

In this case since 2x is same as 2x, it would be \[\cos(2x+2x)=\cos (2x) \cos(2x)- \sin (2x) \sin(2x)=\cos^2(2x)-\sin^2(2x)\] so so far we've got \[\cos^2(2x)-\sin^2(2x)+\cos(2x)=0\]

OpenStudy (solomonzelman):

\[\cos^2(2x)- \color{blue}{ 1-\cos^2(2x) } +\cos(2x)=0\] blue is a substitute for \[\sin^2x\]

OpenStudy (solomonzelman):

\[1+\cos(2x)=0\]

OpenStudy (solomonzelman):

\[\cos(2x)=-1\] do it from here.

OpenStudy (anonymous):

thank you!!!

OpenStudy (solomonzelman):

You welcome!